# SORT3D: Spatial Object-centric Reasoning Toolbox for Zero-Shot 3D Grounding Using Large Language Models

### v1

@@ III-C Spatial Reasoning Toolbox
With only the relevant objects extracted, the subset of objects and their captions are then fed to an LLM reasoning agent as a list of $n_{o}$ objects, and each object $o_{i}$ is represented by the list of attributes $\{id,name,caption,c_{x},c_{y},c_{z},size\}$, where $id$ is a unique integer identifier for the object, and $(c_{x},c_{y},c_{z})$ are discretized coordinates of the object center. The LLM is provided with a spatial reasoning toolbox and prompted to use chain-of-thought reasoning to pick out the referred object in the referential statement. We parse the selection as the navigation target for the real-world experiment (section III-D).
Directly prompting LLMs for object grounding often results in poor performance due to their limitations in spatial and mathematical reasoning: for example, when querying ”the nightstand to the left of the bed”, the LLM picks the one with the smallest $x$ value. As a result, we create a spatial reasoning toolbox consisting of heuristic search functions that find objects referred to by spatial relations and order them in likelihood of satisfying the relations. We then prompt the LLM to delegate object grounding to sequences of calls to the toolbox, which effectively maps operations that require complex spatial reasoning into sequential reasoning and set union/intersection problems. To specifically handle view-dependent queries, we leverage knowledge of free traversable space to anchor feasible viewpoints. During prompting of the LLM, only one in-context example of spatial reasoning with tool calls to demonstrate their usage is required. The modular architecture of our toolbox allows for easy expansion or modification of provided tools.
III-D Parsing for Action Execution
Finally, we tailor our pipeline for downstream navigation tasks by parsing LLM generated actions to either go near a given waypoint or go between two waypoints based on the target object selected. We first convert the target object location into a navigation waypoint. Given the center point of the selected target object, we calculate the closest point in free-space in the scene traversable by a robot agent and call on action functions with this waypoint.
@@ IV EXPERIMENTAL SETUP
We evaluate our method on two 3D object-referential datasets, ReferIt3D [1] and VLA-3D [17]. Both datasets consist of utterances describing a target object in a Scannet [13] scene using spatial relations. In particular, ReferIt3D is split into Sr3D, which consists of synthetically generated utterances from five relation categories while Nr3D consists of natural language statements collected from humans, with unconstrained methods of describing target objects. Statements are categorized as “Easy”/“Hard” based on the number of “distractor” objects of the same class as the target object in the scene and also “View-Dependent”/“View-Independent”. VLA-3D consists entirely of template-generated statements that are view-independent but contains utterances for a diverse set of 3D scans and enforces every statement to contain a spatial relation. The set of spatial relations is expanded to include eight total relations, including ternary relations (e.g. “between”) and numerical relations (e.g. “second closest”). Utterances may also contain attributes such as color and size if needed to disambiguate the target object from distractors.
IV-A 3D Scene Grounding
We test our model on both ReferIt3D subsets and the subset of VLA-3D using Scannet scenes and compare to SOTA baselines. On each data subset, we evaluate our model on 200 sampled statements  [^2] We use a subset of the test set due to the cost of LLM evaluation on the full test set and the need for multiple runs to obtain variance statistics, sampled to match the distribution of Easy, Hard, View-Dependent, and View-Independent statements in the original ReferIt3D test dataset. We focus our comparison against Transcrib3d [6] which, to the best of our knowledge, is the best performing model on ReferIt3D to date. We note, however, that their evaluation skipped over statements in the test set that either do not explicitly mention the target object or human annotators were unable to guess correctly during the dataset generation. We provide results for our method on both cases, with these statements included and skipped, where “+skip” indicates the exclusion of such statements and subsequent simplification of the problem. For a fair comparison, we run Transcrib3d with the same two LLMs we use on the same test splits 333the GPT4 model used in their work is now a legacy model. We run their method with GPT4o instead.. For our methods, we conduct multiple trials on each data split to measure variance in LLMs, reported with standard deviation values on the grounding accuracy, which we note that other LLM-based methods do not report.
@@ V RESULTS AND DISCUSSION
V-A 3D Scene Grounding
The grounding accuracy on ReferIt3D is shown in Tables I and II, and accuracy on VLA-3D is shown in Table III. We see that our method achieves higher accuracy with GPT-4o as the LLM backend and is on par with SOTA methods on View-Dependent statements in Nr3D and Hard statements in VLA-3D while requiring no data to train. We also note that the use of LLMs introduces variance between trials, affecting grounding accuracy up to 6%.
While supervised baselines such as ViL3DRel[28], 3D-VisTA [5], and SceneVerse [19] report slightly higher overall grounding accuracies, these methods are explicitly trained on ReferIt3D data. Similarly, while Transcrib3D [6] reports higher accuracies, it relies on a) removing statements that annotators incorrectly answered and b) iterative self-correction, where incorrect queries are fed back with oracle labels as few-shot examples for fine-tuning. While this approach improves accuracy on the benchmark, it arguably ”forces” the LLM to remember its prior mistakes within the dataset.
In contrast, our approach is purely zero-shot, requiring only one single example of how to use the spatial reasoning toolbox, which does not have to be from a particular dataset, and we employed no dataset-specific training or fine-tuning. Despite this, by mapping spatial reasoning into sequential reasoning, our spatial reasoning toolbox approach achieves performance comparable to SOTA supervised and fine-tuned methods on view-dependent statements in Nr3D, and surpasses other zero-shot methods on grounding accuracy, particularly on view-dependent statements. On Sr3D, SORT3D surpasses SOTA supervised training methods while achieving closer performance to fine-tuned methods. This demonstrates the effectiveness of our approach at handling spatial reasoning where viewpoint anchoring is required. We see that SORT3D is able to explainably resolve complex view-dependent relations with multiple anchors and complex semantic descriptions (Fig. 4-a), and are additionally able to explain model failure points by analyzing its chain of thought (Fig. 4-b).
On VLA-3D, our method surpasses other methods on Hard statements by a large margin. VLA-3D contains a large set of statements using size and color descriptions when referring to objects, which our method effectively utilizes through object filtering, 2D caption descriptions, and the spatial reasoning toolbox to identify target objects with multiple distractors in the scene. Transcrib3D [6], when not fine-tuned toward the dataset’s hard statements, failed to generalize well despite using the same LLM backbone. The results on VLA-3D strongly support the benefits of incorporating 2D captions, as they provide critical, fine-grained object-level attributes that refine the referential grounding.
@@
TABLE I: Performance on the Nr3D dataset from the ReferIt3D benchmark. Asterisks (*) indicate results reported from the paper directly. “+skip” indicates exclusion of statements answered incorrectly by human annotators. “View Dep.” and “View Ind.” stand for view-dependent and view-independent respectively.
@@
Fine-Tuned Methods
Transcrib3D* [6] (GPT-4+skip) 70.2 60.1 75.4
Transcrib3D (GPT-4o) 58.5 54.0 60.6
Transcrib3D (Mistral) 58.5 52.4 61.3
Transcrib3D (GPT-4o+skip) 65.6 63.3 66.7
Transcrib3D (Mistral+skip) 63.8 57.1 66.7
@@
ZSVG3D [20] 39.0 36.8 40.0
VLM-Grounder [11] 48.0 45.8 49.4
CSVG [9] 59.2 53.0 62.5
Ours (Mistral) 56.5$\pm$0.5 57.6$\pm$3.6 56.2$\pm$2.3
Ours (GPT-4o) 60.5$\pm$1.0 56.6$\pm$3.3 62.3$\pm$2.4
Ours (GPT-4o+skip) 62.0$\pm$1.2 56.6$\pm$0.0 64.3$\pm$1.7
Ours (Mistral+skip) 61.6$\pm$0.3 59.4$\pm$0.9 62.6$\pm$0.9
TABLE II: Performance on the Sr3D dataset from the ReferIt3D benchmark. Asterisks (*) indicate results reported from the paper directly.
@@
Fine-Tuned Methods
Transcrib3D* [6] (GPT-4+skip) 98.4 98.2 98.4
@@
Zero-Shot Methods
Ours (Mistral) 92.0$\pm$0.7 90.9$\pm$0.0 92.2$\pm$0.8
Ours (GPT-4o) 92.0$\pm$0.0 95.5$\pm$0.0 91.6$\pm$0.0
TABLE III: Grounding accuracy on VLA-3D test subset
VLA3D
@@ V-B Ablation of Captioning Module
We evaluate the effect on grounding accuracy of adding open-vocabulary captions generated from 2D images of objects in the scene. We augment the Transcrib3D [6] baseline model with our captions for each object as additional information passed into the LLM reasoner. We hypothesize that these finer descriptions of object attributes will help the model in disambiguating objects when given free-form referential statements. We test this hypothesis through an ablation study using GPT-4o with both our method and Transcrib3D on Nr3D, shown in Table IV. In both Transcrib3D and our method, we observe consistent and significant improvements across all statements types. For our approach specifically, the addition of captions improves performance the most (11%) on view-dependent statements. These results demonstrate that understanding detailed object attributes is important for effective 3D grounding and leveraging 2D VLMs is an effective method for this. Many referential statements rely on subtle distinctions—such as color, shape, texture, or affordance—that traditional 3D models often miss.
V-C Real-World Validation
To validate SORT3D’s pipeline in the real-world, we implemented the full system on a differential-drive ground vehicle equipped with a 360-degree camera, a Velodyne LiDAR sensor, an onboard Intel NUC for low level autonomy, and a desktop with an RTX 4090. The system starts with an empty map and continuously integrates real-time RGB-D input to perform panoptic segmentation. This enables the robot to simultaneously build a map, detect objects, and assign open-vocabulary labels, allowing it to navigate and understand its environment dynamically. As the robot explores, it captures images of objects and our pipeline selects image crops and generates open-vocabulary captions, then responds to grounding language queries whenever they are typed in real-time, triggering the rest of the SORT3D pipeline. We validate our system in a previously unseen environment, demonstrating its ability to handle complex statements including compound statements and open-vocabulary object-centric navigation commands, as shown in Fig 6.
Figure 5: An overhead view of the real world kitchen environment and the differential-drive robot used in the experiment.
Figure 6: Real-world examples of SORT3D referential grounding and navigation on compound relational statements and multi-object navigation 3D. The top row shows the egocentric camera view of the robot and the bottom row shows the path taken in a map of the scene. The target objects are highlighted in green and the anchor objects blue.
VI LIMITATIONS AND FUTURE WORK
We acknowledge that limitations exist within in our approach. First, our system relies on internet access to call online APIs, which is a reasonable assumption for indoor environments like homes but may not hold for generalizable robot navigation systems. To address this limitation, the modular design of our pipeline allows for the LLMs to be easily replaced by smaller local models.
We also recognize that our evaluation set is limited. There are few existing benchmarks designed to rigorously test online 3D referential grounding with diverse, attribute-rich natural language. This constraint makes it challenging to fully assess the generalizability of our approach across different environments or while deployed on a system. Evaluating our method in a simulated environment with real-time interactions would provide deeper insights into its effectiveness and adaptability as a system.
TABLE IV: Grounding accuracy with and without captions on Nr3D

### v2

@@ III-D Spatial Reasoning Toolbox
With only the relevant objects extracted, the subset of objects and their captions are then fed to an LLM reasoning agent as a list of $n_{o}$ objects, where each object $o_{i}$ is represented by the list of attributes: $\{\texttt{id},\texttt{name},\texttt{caption},c_{x},c_{y},c_{z},\texttt{size}\}$, where id is a unique integer identifier for the object, $(c_{x},c_{y},c_{z})$ are discretized coordinates of the object center, and size is the area of the largest face.
However, directly prompting off-the-shelf LLMs for 3D referential grounding with this scene representation results in poor performance due to their limitations in spatial and mathematical reasoning: for example, when querying ”the nightstand to the left of the bed”, the LLM picks the one with the smallest $x$ value. As a result, we abstract spatial reasoning away from the LLM by creating a spatial reasoning toolbox consisting of heuristic search functions that find objects referred to by a set of elementary spatial relations. The key idea is that when referring to an object using relative spatial relations, only a fixed set of relations are needed. Thus, defining this set of heuristic functions is sufficient for handling diverse inter-object referential scenarios. The search functions and their arguments are listed in Table I.
Spatial Search Functions
find_near(target, anchor)*
find_between(target, anchor1, anchor2)*
find_above(target, anchor)*
find_below(target, anchor)*
order_bottom_to_top(targets)*
order_smallest_to_largest(targets)*
find_objects_near_room_corner(targets)*
find_left(target, anchor)
find_right(target, anchor)
order_left_to_right(target, anchor)
TABLE I: Heuristic search functions in the spatial toolbox. View-independent functions are marked with an asterisk (*), and view-dependent functions are marked with a dagger (†).
The LLM is prompted with an in-context example to decompose a referential statement into a series of search calls and choose a single object ID from the returned lists of IDs. For example, given the query “Find the computer near the desk with a printer on it”, the LLM first calls find_below(desk, printer) returning [2], then find_near(computer, 2) returning [3, 4], finally picking the computer with ID 3. We note a few implementation details regarding the search functions:
  1. 1.
    For statements employing view-dependent relationships, like left and right, we tackle the case where no specific observer viewpoint is given (common in Nr3D), and assume the relationship is unambiguous from any feasible viewing direction. We therefore determinine left/right direction from the signed angle formed by the nearest point in free space and the target and anchor’s centroids.
  2. 2.
    We use the area of the largest face in order_smallest_to_largest, which is more intuitive than volume for flat objects.
  3. 3.
    All search functions have access to the full output of the perception module including object names and 3D bounding boxes, which is partially hidden from the LLM. This allows us to develop more complex logic for search functions without affecting LLM reasoning, which is an advantage of our approach.
Our accompanying repository contains further implementation details on the spatial search functions.
III-E Parsing for Action Execution
For our benchmark results, the LLM is prompted to only output a single object ID denoting the chosen object. For our real-world deployed pipeline, the LLM may output a series of either go_near or go_between function calls which take object IDs as arguments and generate waypoints in the scene which are sent to an obstacle-avoidant downstream planner for sequential navigation. go_near(target) places a waypoint at the closest point in traversable space to the target’s center, while go_between(target1, target2) places one near the midpoint between the two targets.
@@ IV Experimental Setup
We quantitatively evaluate our method on two 3D object-referential datasets, ReferIt3D [1] and IRef-VLA [17]. Both datasets consist of utterances describing a target object in a ScanNet [13] scene using spatial relations. In particular, ReferIt3D is split into Sr3D, which consists of synthetically generated utterances from five relation categories while Nr3D consists of natural language statements collected from humans, with unconstrained methods of describing target objects. Statements are categorized as “Easy”/“Hard” based on the number of “distractor” objects of the same class as the target object in the scene and also “View-Dependent”/“View-Independent”. IRef-VLA consists entirely of template-generated statements that are view-independent but contains utterances for a diverse set of 3D scans and enforces every statement to contain a spatial relation. The set of spatial relations is expanded to include eight total relations, including ternary relations (e.g. “between”) and numerical relations (e.g. “second closest”). Utterances may also contain attributes such as color and size if needed to disambiguate the target object from distractors. Additionally, we deploy SORT3D on two ground vehicles, and validate the system’s generalizability by testing navigation commands containing references to spatial relations, references to object attributes, and implicit or indirect requests in three previously unseen indoor environments.
IV-A Referential Grounding on Benchmark Datasets
We test our model on both ReferIt3D subsets and the subset of IRef-VLA using ScanNet scenes and compare to SOTA baselines. On each data subset, we evaluate our model on 200 sampled statements [^3] We use a subset of the test set due to the cost of LLM evaluation on the full test set and the need for multiple runs to obtain variance statistics, sampled to match the distribution of Easy, Hard, View-Dependent, and View-Independent statements in the original ReferIt3D test dataset. We focus our comparison against Transcrib3D [6] which, to the best of our knowledge, is the best performing model on ReferIt3D to date. For a fair comparison, we run Transcrib3D with the same two LLMs we use on the same test splits444the GPT4 model used in their work is now a legacy model. We run their method with GPT4o instead.. For our methods, we conduct multiple trials on each data split to measure variance in LLMs, reported with standard deviation values on the grounding accuracy, which we note that other LLM-based methods do not report.
@@ V Results and Discussion
V-A Referential Grounding on Benchmark Datasets
The grounding accuracy on ReferIt3D is shown in Tables II and III, and accuracy on IRef-VLA is shown in Table IV. We see that our method achieves higher accuracy with GPT-4o as the LLM backend and is on par with SOTA methods on View-Dependent statements in Nr3D and Hard statements in IRef-VLA while requiring no data to train. We also note that the use of LLMs introduces variance between trials, affecting grounding accuracy up to 6%.
While supervised baselines such as ViL3DRel [20], 3D-VisTA [5], and SceneVerse [19] report slightly higher overall grounding accuracies, these methods are explicitly trained on ReferIt3D data. Similarly, while Transcrib3D reports higher accuracies, it relies on guiding principles [6] that are tailored to the language used in Nr3D and Sr3D, which improve performance on those benchmarks.
In contrast, our approach is purely zero-shot, requiring only one single example of how to use the spatial reasoning toolbox, which does not have to be from a particular dataset, and we employed no dataset-specific training or fine-tuning. Despite this, by leveraging foundation models to obtain object semantic attributes and mapping spatial reasoning into sequential reasoning, our spatial reasoning toolbox approach achieves overall performance comparable to SOTA supervised and fine-tuned methods on Nr3D, and performance on view-dependent statements on par with Transcrib3D. On Sr3D, SORT3D surpasses SOTA supervised training methods and achieves close overall performance to Transcrib3D while surpassing it in view-dependent accuracy. This demonstrates the effectiveness of our approach at handling spatial reasoning where viewpoint anchoring is required. We see that SORT3D is able to explainably resolve complex view-dependent relations with multiple anchors and complex semantic descriptions (Figure 4-a), while also providing explainable model failure points by analyzing its chain of thought reasoning (Figure 4-b).
On IRef-VLA, our method surpasses other methods on Hard statements by a large margin. IRef-VLA contains a large set of statements using size and color descriptions when referring to objects, which our method effectively grounds by utilizing object filtering, semantic attributes captured by captions, and the spatial reasoning toolbox to identify target objects with multiple distractors in the scene. Transcrib3D, when used in its zero-shot formulation with principles and not fine-tuned on the dataset’s hard statements [6], fails to generalize well despite using the same LLM backbone. The results on IRef-VLA strongly support the benefits of incorporating 2D captions, as they provide critical, fine-grained object-level attributes that refine the referential grounding.
@@ V-A Referential Grounding on Benchmark Datasets
TABLE II: Performance on the Nr3D dataset from the ReferIt3D benchmark. Asterisks (*) indicate results reported from the paper directly. “View Dep.” and “View Ind.” stand for view-dependent and view-independent respectively.
@@
ZSVG3D* [21] 39.0 36.8 40.0
VLM-Grounder* [11] 48.0 45.8 49.4
CSVG* [9] 59.2 53.0 62.5
Transcrib3D* [6] (GPT-4) 70.2 60.1 75.4
Transcrib3D (GPT-4o) 65.6 63.3 66.7
Transcrib3D (Mistral) 63.8 57.1 66.7
[…]
TABLE III: Performance on the Sr3D dataset from the ReferIt3D benchmark. Asterisks (*) indicate results reported from the paper directly.
@@
Zero-Shot Methods
Transcrib3D* [6] (GPT-4) 98.4 98.2 98.4
@@
[…]
TABLE IV: Grounding accuracy on IRef-VLA test subset
IRef-VLA
@@ V-B Ablation of Captioning Module
We evaluate the effect on grounding accuracy of adding open-vocabulary captions generated from 2D images of objects in the scene. We augment the Transcrib3D [6] baseline model with our captions for each object as additional information passed into the LLM reasoner. We hypothesize that these finer descriptions of object attributes will help the model in disambiguating objects when given free-form referential statements. We test this hypothesis through an ablation study using GPT-4o with both our method and Transcrib3D on Nr3D, shown in Table V. In both Transcrib3D and our method, we observe consistent and significant improvements across all statements types. For our approach specifically, the addition of captions improves performance the most (11%) on view-dependent statements. These results demonstrate that understanding detailed object attributes is important for effective 3D grounding and leveraging 2D VLMs is an effective method for this. Many referential statements rely on subtle distinctions—such as color, shape, texture, or affordance—that traditional 3D models often miss.
TABLE V: Grounding accuracy with and without captions on Nr3D
@@
V-C Real-World Validation
To validate SORT3D’s pipeline in the real-world, we implement the full system on two autonomous mobile robots: a mecanum-wheeled robot and a differential drive robot with a wheelchair base, each equipped with 360-degree cameras, respective Livox and Velodyne LiDARS, onboard Intel NUCs for low level autonomy, and RTX 4090s with 16GB and 24GB of VRAM respectively. We validate our system in two previously unseen indoor environments: a student lounge (Figure 6, and a university corridor (Figure 5). We attempt two different types of navigational queries, which target the system’s ability to ground statements that employ both spatial references and object semantic attributes. Before issuing a query, we navigate each robot around the scene to build a semantic map after prompting the open vocabulary detector with the names and synonyms of objects in the scene (shown in RViz in the bottom of Figures 6 and 5), and collect image crops for each detected object. As an implementation detail, captions are batch-generated only after a user query is typed into the system to speed up semantic mapping and decrease overall power consumption. RViz visualizations of the instance level semantic maps, the objects and corresponding waypoints chosen by the grounding model for each statement, and pictures of each platform navigating through its environment are shown in Figures 6 and 5. In each statement we test, SORT3D successfully grounds one or more referenced objects, demonstrating the versatility of our approach for grounding complex expressions involving spatial references and semantic attributes in previously unseen scenes. Further experiments on more environments and types of statements are found in our accompanying repository and videos.
Figure 5: Navigation on the mecanum robot in the university corridor given the statement “Go to the table next to the bookshelf, then to the chair next to the plant.” The system successfully navigates to both spatially referenced objects.
Figure 6: Navigation on the mecanum robot in the student lounge given the statement “I want to play a board game, fetch me one from the shelf.” The system successfully navigates to the shelf with the board game on it by referring to semantic caption descriptions.
VI Limitations and Future Work
We acknowledge that limitations exist within our approach. First, our system relies on internet access to call online APIs, which is a reasonable assumption for indoor environments like homes and offices, but may not hold for other environments where generalizable robot systems may need to be deployed. To address this limitation, the modular design of our pipeline allows for the LLMs to be easily replaced by smaller local models.
We also recognize that our evaluation set is limited. There are few existing benchmarks designed to rigorously test online 3D referential grounding with diverse, attribute-rich natural language. This constraint makes it challenging to fully assess the generalizability of our approach across different environments or while deployed on a system. Evaluating our method in a simulated environment with real-time interactions would provide deeper insights into its effectiveness and adaptability as a system. It would also allow us to investigate more practical settings such as multi-turn grounding scenarios and failure correction.